Resource Supply and Usage, Special Order, Relevancy

Elliott, Inc., has four salaried clerks to process purchase orders. Each clerk is paid a salary of $26,500 and is capable of processing as many as 6,550 purchase orders per year. Each clerk uses a PC and laser printer in processing orders. Time available on each PC system is sufficient to process 6,550 orders per year. The cost of each PC system is $1,150 per year. In addition to the salaries, Elliott spends $26,724 for forms, postage, and other supplies (assuming 26,200 purchase orders are processed). During the year, 25,400 orders were processed.

Required:

1. Which of the resources associated with purchasing can be classified as flexible resources?

Forms and clerks

Which of the resources associated with purchasing can be classified as committed resources?

Postage and PC system

2. Compute the total activity availability, and break this into activity usage and unused activity.

Activity availability: purchase orders
Activity usage: purchase orders
Unused activity: purchase orders

3. Calculate the total cost of resources supplied (activity cost), and break this into the cost of activity used and the cost of unused activity. When required, round your answers to the nearest dollar. If required, round your intermediate calculations to five decimal places.

Activity cost: $
Cost of activity used: $
Cost of unused activity: $
Answer #1

ANSWER

1

Forms and clerks

Postage and PC system

2] Activity availability 26,200 purchase orders
Activity usage 25,400 purchase orders
Unused activity 8,00 purchase orders
3] Activity cost =26,500 *4+1,150 *4+(26,724 /26,200 )*25,400 = $          136,508
Cost of activity used = (26,500 *4+1,150 *4)*25,400 /26,200 +(26,724 /26,200 )*25,400 = $          133,124
Cost of unused activity = (26,500 *4+1,150 *4)*(26,200 -25,400)/26,200 = $ 3,377

_____________________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************
Answer #2

ANSWER

1

Forms and clerks

Postage and PC system

2] Activity availability 26,200 purchase orders
Activity usage 25,400 purchase orders
Unused activity 8,00 purchase orders
3] Activity cost =26,500 *4+1,150 *4+(26,724 /26,200 )*25,400 = $          136,508
Cost of activity used = (26,500 *4+1,150 *4)*25,400 /26,200 +(26,724 /26,200 )*25,400 = $          133,124
Cost of unused activity = (26,500 *4+1,150 *4)*(26,200 -25,400)/26,200 = $ 3,377

_____________________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************